# A Cross-Scene Benchmark for Open-World Drone Active Tracking

### v1

@@ Appendix D Baselines
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
AOT $49_{\pm 3}$ $49_{\pm 9}$ $45_{\pm 5}$ $49_{\pm 3}$ $48_{\pm 3}$ $48_{\pm 3}$ $49_{\pm 4}$ $49_{\pm 3}$ $49_{\pm 3}$
R-VAT $\textbf{279}_{\pm 110}$ $\textbf{129}_{\pm 112}$ $\textbf{153}_{\pm 119}$ $\textbf{135}_{\pm 109}$ $\textbf{112}_{\pm 92}$ $\textbf{191}_{\pm 122}$ $\textbf{257}_{\pm 126}$ $\textbf{316}_{\pm 84}$ $\textbf{202}_{\pm 119}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
AOT $9_{\pm 0}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 0}$ $9_{\pm 0}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$
R-VAT $\textbf{278}_{\pm 111}$ $\textbf{307}_{\pm 124}$ $\textbf{305}_{\pm 94}$ $\textbf{119}_{\pm 110}$ $\textbf{170}_{\pm 139}$ $\textbf{275}_{\pm 121}$ $\textbf{182}_{\pm 131}$ $\textbf{307}_{\pm 124}$ $\textbf{307}_{\pm 97}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
D-VAT $46_{\pm 8}$ $45_{\pm 9}$ $44_{\pm 8}$ $69_{\pm 42}$ $45_{\pm 8}$ $45_{\pm 3}$ $44_{\pm 8}$ $44_{\pm 8}$ $43_{\pm 8}$
R-VAT $\textbf{234}_{\pm 122}$ $\textbf{160}_{\pm 139}$ $\textbf{239}_{\pm 134}$ $\textbf{93}_{\pm 102}$ $\textbf{153}_{\pm 115}$ $\textbf{140}_{\pm 118}$ $\textbf{257}_{\pm 122}$ $\textbf{257}_{\pm 120}$ $\textbf{114}_{\pm 115}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
D-VAT $8_{\pm 1}$ $8_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 1}$ $8_{\pm 1}$ $8_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 2}$
R-VAT $\textbf{209}_{\pm 131}$ $\textbf{184}_{\pm 136}$ $\textbf{202}_{\pm 129}$ $\textbf{203}_{\pm 119}$ $\textbf{189}_{\pm 93}$ $\textbf{223}_{\pm 114}$ $\textbf{167}_{\pm 135}$ $\textbf{165}_{\pm 126}$ $\textbf{178}_{\pm 125}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{112}_{\pm 86}$ $\textbf{144}_{\pm 110}$ $\textbf{203}_{\pm 133}$ $\textbf{143}_{\pm 134}$ $\textbf{181}_{\pm 116}$ $\textbf{214}_{\pm 111}$ $\textbf{190}_{\pm 129}$ $\textbf{168}_{\pm 110}$ $\textbf{99}_{\pm 67}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{162}_{\pm 89}$ $\textbf{170}_{\pm 125}$ $\textbf{237}_{\pm 128}$ $\textbf{81}_{\pm 71}$ $\textbf{159}_{\pm 119}$ $\textbf{243}_{\pm 117}$ $\textbf{253}_{\pm 109}$ $\textbf{245}_{\pm 117}$ $\textbf{168}_{\pm 105}$
Table 12: The detailed results of comparison experiments on $CR$ metric. Each row in the table represents the training results on a specific map, e.g., the first row corresponds to the results trained on the map citystreet-day. Each column represents the testing results on the corresponding map, e.g., the second column corresponds to the testing results on the map desert-day.
Figure 14: Schematic diagram of reward curves on DAT scenes. The red curve represents the reward of the proposed R-VAT, while the yellow curve corresponds to the R-VAT (w/o CBT).
@@
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
AOT $0.25_{\pm 0.02}$ $0.24_{\pm 0.03}$ $0.22_{\pm 0.03}$ $0.25_{\pm 0.02}$ $0.23_{\pm 0.03}$ $0.24_{\pm 0.01}$ $0.25_{\pm 0.02}$ $0.25_{\pm 0.02}$ $0.24_{\pm 0.02}$
R-VAT $\textbf{0.80}_{\pm 0.30}$ $\textbf{0.54}_{\pm 0.32}$ $\textbf{0.50}_{\pm 0.32}$ $\textbf{0.45}_{\pm 0.30}$ $\textbf{0.44}_{\pm 0.24}$ $\textbf{0.66}_{\pm 0.27}$ $\textbf{0.72}_{\pm 0.29}$ $\textbf{0.93}_{\pm 0.14}$ $\textbf{0.79}_{\pm 0.24}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
AOT $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.01}$
R-VAT $\textbf{0.73}_{\pm 0.31}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{0.87}_{\pm 0.19}$ $\textbf{0.38}_{\pm 0.32}$ $\textbf{0.56}_{\pm 0.28}$ $\textbf{0.82}_{\pm 0.25}$ $\textbf{0.57}_{\pm 0.31}$ $\textbf{0.86}_{\pm 0.28}$ $\textbf{0.86}_{\pm 0.22}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
D-VAT $0.23_{\pm 0.04}$ $0.23_{\pm 0.04}$ $0.22_{\pm 0.05}$ $0.31_{\pm 0.14}$ $0.24_{\pm 0.06}$ $0.22_{\pm 0.01}$ $0.22_{\pm 0.04}$ $0.22_{\pm 0.05}$ $0.23_{\pm 0.05}$
R-VAT $\textbf{0.72}_{\pm 0.28}$ $\textbf{0.51}_{\pm 0.34}$ $\textbf{0.73}_{\pm 0.32}$ $\textbf{0.46}_{\pm 0.29}$ $\textbf{0.59}_{\pm 0.33}$ $\textbf{0.48}_{\pm 0.31}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.40}_{\pm 0.29}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
D-VAT $0.05_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.77}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.67}_{\pm 0.29}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.49}_{\pm 0.29}$ $\textbf{0.63}_{\pm 0.33}$ $\textbf{0.58}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.29}$ $\textbf{0.64}_{\pm 0.28}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
D-VAT $0.25_{\pm 0.04}$ $0.23_{\pm 0.04}$ $0.23_{\pm 0.05}$ $0.30_{\pm 0.15}$ $0.26_{\pm 0.06}$ $0.22_{\pm 0.01}$ $0.26_{\pm 0.06}$ $0.26_{\pm 0.06}$ $0.25_{\pm 0.06}$
R-VAT $\textbf{0.43}_{\pm 0.25}$ $\textbf{0.47}_{\pm 0.30}$ $\textbf{0.64}_{\pm 0.31}$ $\textbf{0.43}_{\pm 0.28}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{0.59}_{\pm 0.30}$ $\textbf{0.59}_{\pm 0.39}$ $\textbf{0.62}_{\pm 0.32}$ $\textbf{0.41}_{\pm 0.24}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
AOT $0.24_{\pm 0.02}$ $0.24_{\pm 0.04}$ $0.22_{\pm 0.03}$ $0.25_{\pm 0.02}$ $0.24_{\pm 0.02}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$ $0.23_{\pm 0.01}$
D-VAT $0.07_{\pm 0.01}$ $0.07_{\pm 0.01}$ $0.07_{\pm 0.00}$ $0.08_{\pm 0.01}$ $0.07_{\pm 0.00}$ $0.07_{\pm 0.00}$ $0.08_{\pm 0.00}$ $0.07_{\pm 0.00}$ $0.08_{\pm 0.00}$
R-VAT $\textbf{0.48}_{\pm 0.24}$ $\textbf{0.59}_{\pm 0.34}$ $\textbf{0.72}_{\pm 0.26}$ $\textbf{0.33}_{\pm 0.20}$ $\textbf{0.58}_{\pm 0.28}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{0.67}_{\pm 0.32}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.51}_{\pm 0.28}$
Table 13: The detailed results of comparison experiments on $TSR$ metric. Each row in the table represents the training results on a specific map, e.g., the second row corresponds to the results trained on the map desert-day. Each column represents the testing results on the corresponding map, e.g., the second column corresponds to the testing results on the map desert-day.
@@ Appendix E More Experiments
E.1 Settings
@@
Training Steps. Due to the varying challenges posed by different scene maps, the convergence speed of the agent differs across experiments, resulting in a unique total training step count for each map. The number of training steps for the agent across the six scenarios is shown in Tab. 10.
@@
Ablation Experiment Settings. In this section, we introduce the training conditions of the single-stage reinforcement learning method and R-VAT, as well as the criteria for stage transitions. In single-stage reinforcement learning, the agent is placed in one of six scenarios (citystreet, desert, village, downtown, lake, and farmland) for training. For R-VAT, the agent is first trained in an environment where a randomly colored target moves straight along a line without obstacles. After convergence, the model is then trained in the corresponding complex scenarios. Specifically, the transition steps T for R-VAT are shown in Tab. 11.
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{279}_{\pm 110}$ $\textbf{129}_{\pm 112}$ $\textbf{153}_{\pm 119}$ $\textbf{135}_{\pm 109}$ $\textbf{112}_{\pm 92}$ $\textbf{191}_{\pm 122}$ $\textbf{257}_{\pm 126}$ $\textbf{316}_{\pm 84}$ $\textbf{202}_{\pm 119}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{278}_{\pm 111}$ $\textbf{307}_{\pm 124}$ $\textbf{305}_{\pm 94}$ $119_{\pm 110}$ $170_{\pm 139}$ $\textbf{275}_{\pm 121}$ $182_{\pm 131}$ $\textbf{307}_{\pm 124}$ $\textbf{307}_{\pm 97}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{234}_{\pm 122}$ $160_{\pm 139}$ $239_{\pm 134}$ $\textbf{93}_{\pm 102}$ $\textbf{153}_{\pm 115}$ $140_{\pm 118}$ $257_{\pm 122}$ $\textbf{257}_{\pm 120}$ $\textbf{114}_{\pm 115}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{209}_{\pm 131}$ $\textbf{184}_{\pm 136}$ $\textbf{202}_{\pm 129}$ $\textbf{203}_{\pm 119}$ $\textbf{189}_{\pm 93}$ $\textbf{223}_{\pm 114}$ $\textbf{167}_{\pm 135}$ $\textbf{165}_{\pm 126}$ $\textbf{178}_{\pm 125}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $112_{\pm 86}$ $\textbf{144}_{\pm 110}$ $\textbf{203}_{\pm 133}$ $\textbf{143}_{\pm 134}$ $181_{\pm 116}$ $\textbf{214}_{\pm 111}$ $\textbf{190}_{\pm 129}$ $168_{\pm 110}$ $99_{\pm 67}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{162}_{\pm 89}$ $\textbf{170}_{\pm 125}$ $\textbf{237}_{\pm 128}$ $\textbf{81}_{\pm 71}$ $\textbf{159}_{\pm 119}$ $\textbf{243}_{\pm 117}$ $\textbf{253}_{\pm 109}$ $\textbf{245}_{\pm 117}$ $\textbf{168}_{\pm 105}$
Table 14: Effectiveness of Curriculum-Based Training strategy on the DAT benchmark, results from $CR$ metric.
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{0.80}_{\pm 0.30}$ $\textbf{0.54}_{\pm 0.32}$ $\textbf{0.50}_{\pm 0.32}$ $\textbf{0.45}_{\pm 0.30}$ $\textbf{0.44}_{\pm 0.24}$ $\textbf{0.66}_{\pm 0.27}$ $\textbf{0.72}_{\pm 0.29}$ $\textbf{0.93}_{\pm 0.14}$ $\textbf{0.79}_{\pm 0.24}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $0.73_{\pm 0.31}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{0.87}_{\pm 0.19}$ $0.38_{\pm 0.32}$ $0.56_{\pm 0.28}$ $\textbf{0.82}_{\pm 0.25}$ $0.57_{\pm 0.31}$ $\textbf{0.86}_{\pm 0.28}$ $\textbf{0.86}_{\pm 0.22}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $0.72_{\pm 0.28}$ $0.51_{\pm 0.34}$ $0.73_{\pm 0.32}$ $\textbf{0.46}_{\pm 0.29}$ $\textbf{0.59}_{\pm 0.33}$ $0.48_{\pm 0.31}$ $0.71_{\pm 0.32}$ $\textbf{0.71}_{\pm 0.32}$ $0.40_{\pm 0.29}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $0.29_{\pm 0.04}$ $0.27_{\pm 0.03}$ $0.27_{\pm 0.03}$ $0.33_{\pm 0.06}$ $0.28_{\pm 0.03}$ $0.27_{\pm 0.01}$ $0.33_{\pm 0.06}$ $0.33_{\pm 0.06}$ $0.33_{\pm 0.06}$
R-VAT $\textbf{0.77}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.67}_{\pm 0.29}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.49}_{\pm 0.29}$ $\textbf{0.63}_{\pm 0.33}$ $\textbf{0.58}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.29}$ $\textbf{0.64}_{\pm 0.28}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
w/o CBT $\textbf{0.51}_{\pm 0.30}$ $\textbf{0.47}_{\pm 0.29}$ $0.45_{\pm 0.22}$ $\textbf{0.44}_{\pm 0.23}$ $0.57_{\pm 0.28}$ $0.59_{\pm 0.26}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.62}_{\pm 0.24}$ $0.33_{\pm 0.15}$
R-VAT $0.43_{\pm 0.25}$ $0.47_{\pm 0.30}$ $\textbf{0.64}_{\pm 0.31}$ $0.43_{\pm 0.28}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{0.59}_{\pm 0.30}$ $0.59_{\pm 0.39}$ $0.62_{\pm 0.32}$ $\textbf{0.41}_{\pm 0.24}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{0.48}_{\pm 0.24}$ $\textbf{0.59}_{\pm 0.34}$ $\textbf{0.72}_{\pm 0.26}$ $\textbf{0.33}_{\pm 0.20}$ $\textbf{0.58}_{\pm 0.28}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{0.67}_{\pm 0.32}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.51}_{\pm 0.28}$
Table 15: Effectiveness of Curriculum-Based Training strategy on the DAT benchmark, results from $TSR$ metric.
@@ E.2 Comparison Experiments
We provide a comprehensive analysis of the comparative experimental results. Specifically, we provide detailed evaluations for within-scene testing (same scenes, same weather), cross-scene testing (same weather, different scenes) and cross-domain testing (same scene, different weather). Tab. 12 reports the $CR$ metric of three models under cross-scene and cross-domain conditions, while Tab. 13 presents the $TSR$ metric.
@@
As shown in Tab. 12 and Tab. 13, the proposed R-VAT significantly outperforms existing state-of-the-art methods. Due to the reward design based on physical distance, both the AOT [33] and D-VAT [18] fail to accurately reflect the agent’s tracking performance from a top-down perspective (see Sec. C.1 for theoretical proof), leading to misleading training signals for the tracker. Consequently, neither AOT nor D-VAT can effectively learn meaningful features, resulting in irregular performance distributions. In contrast, the proposed R-VAT achieves superior convergence across all scenes. Specifically, in cross-scene experiments, the testing performance of the agent on the downtown map is relatively low, indicating that dense buildings and complex road elements pose significant challenges to the agent. Conversely, the testing performance on the village map is comparatively high, suggesting that the uniform color and simpler road conditions in the village map present fewer challenges.
@@
For cross-domain testing experiments, the agent performs well under night and foggy conditions but struggles under snow conditions. This indicates that the proposed R-VAT exhibits strong robustness to changes in lighting and visibility but is less adaptive to variations in scene tone.
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{279}_{\pm 110}$ $\textbf{129}_{\pm 112}$ $\textbf{153}_{\pm 119}$ $\textbf{135}_{\pm 109}$ $\textbf{112}_{\pm 92}$ $\textbf{191}_{\pm 122}$ $\textbf{257}_{\pm 126}$ $\textbf{316}_{\pm 84}$ $\textbf{202}_{\pm 119}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{278}_{\pm 111}$ $\textbf{307}_{\pm 124}$ $\textbf{305}_{\pm 94}$ $\textbf{119}_{\pm 110}$ $\textbf{170}_{\pm 139}$ $\textbf{275}_{\pm 121}$ $\textbf{182}_{\pm 131}$ $\textbf{307}_{\pm 124}$ $\textbf{307}_{\pm 97}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{234}_{\pm 122}$ $\textbf{160}_{\pm 139}$ $\textbf{239}_{\pm 134}$ $\textbf{93}_{\pm 102}$ $\textbf{153}_{\pm 115}$ $\textbf{140}_{\pm 118}$ $\textbf{257}_{\pm 122}$ $\textbf{257}_{\pm 120}$ $\textbf{114}_{\pm 115}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{209}_{\pm 131}$ $\textbf{184}_{\pm 136}$ $\textbf{202}_{\pm 129}$ $\textbf{203}_{\pm 119}$ $\textbf{189}_{\pm 93}$ $\textbf{223}_{\pm 114}$ $\textbf{167}_{\pm 135}$ $\textbf{165}_{\pm 126}$ $\textbf{178}_{\pm 125}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $11_{\pm 3}$ $11_{\pm 1}$ $9_{\pm 1}$ $9_{\pm 2}$ $9_{\pm 0}$ $8_{\pm 0}$ $9_{\pm 0}$ $10_{\pm 1}$ $8_{\pm 1}$
R-VAT $\textbf{112}_{\pm 86}$ $\textbf{144}_{\pm 110}$ $\textbf{203}_{\pm 133}$ $\textbf{143}_{\pm 134}$ $\textbf{181}_{\pm 116}$ $\textbf{214}_{\pm 111}$ $\textbf{190}_{\pm 129}$ $\textbf{168}_{\pm 110}$ $\textbf{99}_{\pm 67}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
R-VAT $\textbf{162}_{\pm 89}$ $\textbf{170}_{\pm 125}$ $\textbf{237}_{\pm 128}$ $\textbf{81}_{\pm 71}$ $\textbf{159}_{\pm 119}$ $\textbf{243}_{\pm 117}$ $\textbf{253}_{\pm 109}$ $\textbf{245}_{\pm 117}$ $\textbf{168}_{\pm 105}$
Table 16: Effectiveness of reward design on the DAT benchmark, results from $CR$ metric.
Within / Cross Scene Cross Domain
Training Scene: citystreet-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.06_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.80}_{\pm 0.30}$ $\textbf{0.54}_{\pm 0.32}$ $\textbf{0.50}_{\pm 0.32}$ $\textbf{0.45}_{\pm 0.30}$ $\textbf{0.44}_{\pm 0.24}$ $\textbf{0.66}_{\pm 0.27}$ $\textbf{0.72}_{\pm 0.29}$ $\textbf{0.93}_{\pm 0.14}$ $\textbf{0.79}_{\pm 0.24}$
Training Scene: desert-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.10_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.09_{\pm 0.01}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.73}_{\pm 0.31}$ $\textbf{0.84}_{\pm 0.29}$ $\textbf{0.87}_{\pm 0.19}$ $\textbf{0.38}_{\pm 0.32}$ $0.56_{\pm 0.28}$ $\textbf{0.82}_{\pm 0.25}$ $\textbf{0.57}_{\pm 0.31}$ $\textbf{0.86}_{\pm 0.28}$ $\textbf{0.86}_{\pm 0.22}$
Training Scene: village-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.72}_{\pm 0.28}$ $\textbf{0.51}_{\pm 0.34}$ $\textbf{0.73}_{\pm 0.32}$ $\textbf{0.46}_{\pm 0.29}$ $\textbf{0.59}_{\pm 0.33}$ $\textbf{0.48}_{\pm 0.31}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.71}_{\pm 0.32}$ $\textbf{0.40}_{\pm 0.29}$
Training Scene: downtown-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.77}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.67}_{\pm 0.29}$ $\textbf{0.65}_{\pm 0.30}$ $\textbf{0.49}_{\pm 0.29}$ $\textbf{0.63}_{\pm 0.33}$ $\textbf{0.58}_{\pm 0.31}$ $\textbf{0.65}_{\pm 0.29}$ $\textbf{0.64}_{\pm 0.28}$
Training Scene: lake-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.10_{\pm 0.01}$ $0.09_{\pm 0.01}$ $0.07_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.08_{\pm 0.00}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.43}_{\pm 0.25}$ $\textbf{0.47}_{\pm 0.30}$ $\textbf{0.64}_{\pm 0.31}$ $\textbf{0.43}_{\pm 0.28}$ $\textbf{0.61}_{\pm 0.31}$ $\textbf{0.59}_{\pm 0.30}$ $\textbf{0.59}_{\pm 0.39}$ $\textbf{0.62}_{\pm 0.32}$ $\textbf{0.41}_{\pm 0.24}$
Training Scene: farmland-day citystreet desert village downtown lake farmland night foggy snow
$\text{R}_{\text{D-VAT}}$ $0.06_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.01}$ $0.05_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$ $0.06_{\pm 0.00}$
R-VAT $\textbf{0.48}_{\pm 0.24}$ $\textbf{0.59}_{\pm 0.34}$ $\textbf{0.72}_{\pm 0.26}$ $\textbf{0.33}_{\pm 0.20}$ $\textbf{0.58}_{\pm 0.28}$ $\textbf{0.68}_{\pm 0.32}$ $\textbf{0.67}_{\pm 0.32}$ $\textbf{0.78}_{\pm 0.22}$ $\textbf{0.51}_{\pm 0.28}$
Table 17: Effectiveness of reward design on the DAT benchmark, results from $CR$ metric. Each row in the table represents the training results on a specific map, e.g., the first row corresponds to the results trained on the map citystreet-day. Each column represents the testing results on the corresponding map, e.g., the second column corresponds to the testing results on the map desert-day.

### v2

@@ Appendix E More Experiments
E.1 Experiment Settings
More Implementation Details. The training involves a range of 9.2M to 21.3M steps across 35 parallel environments. The webots runs at $500\text{Hz}$, with the algorithm updating every four steps ($125\text{Hz}$). Episodes last up to 1500 steps and were terminated early if the drone lost the target for over 100 consecutive steps, collided, or crashed. The drone translation speed is set to $40\text{m/s}$, and rotational speed to $2\text{rad/s}$. The map features 40 vehicles, each with a maximum speed of $20\text{m/s}$ and acceleration of $\pm 25\text{m/}\text{s}^{2}$. During testing, the altitude is set to $22\text{m}$, the pitch angle to $1.37\text{rad}$, and the target initializes at the camera’s center.
The drone’s translation speed is set to a higher value to prevent it from becoming too similar to the target’s speed (with a maximum of 20 m/s). This prevents simple forward movement from yielding excessively high reward evaluations. If the drone’s speed is set lower (e.g., 20 m/s), it may adopt a suboptimal strategy, relying solely on one action.
@@ E.1 Experiment Settings
Due to the varying challenges posed by different scene maps, the convergence speed of the agent differs across experiments. The training steps are shown in Table 10.
@@
Ablation Experiment Settings. In this section, we introduce the training conditions of the single-stage RL and GC-VAT, as well as the criteria for stage transitions. In single-stage RL, the agent is placed in one of six scenarios (citystreet, desert, village, downtown, lake, and farmland) for training. For GC-VAT, the agent is first trained in an environment where a randomly colored target moves straight along a line without obstacles. After convergence, the model is then trained in the corresponding complex scenarios. The transition steps T for GC-VAT are in Table 11.
Figure 15: Qualitative results on the car8 video.
Table 18: Results (metric is Correct Action Rate) of 8 videos in VOT benchmark.
Video car1 car3 car6 car8
Random $0.418$ $0.434$ $0.418$ $0.430$
Ours 0.696 0.845 0.754 0.833
Video carchase car16 following car9
Random $0.429$ $0.421$ $0.314$ $0.439$
Ours 0.870 0.834 0.773 0.756
Figure 16: Qualitative results on the Car4 video.
Table 19: Results of 8 videos in DTB70 benchmark.
Video Car2 Car4 Car5 RcCar4
Random $0.419$ $0.421$ $0.429$ $0.436$
Ours 0.757 0.894 0.893 0.876
Video Car8 RaceCar RaceCar1 RcCar3
Random $0.411$ $0.462$ $0.430$ $0.400$
Ours 0.803 0.713 0.880 0.851
Figure 17: Qualitative results on the RaceCar1 video.
Table 20: Results of 8 videos in UAVDT UAVDT benchmark.
Video S1603 S0201 S0101 S0306
Random $0.435$ $0.438$ $0.422$ $0.437$
Ours 0.896 0.806 0.865 0.773
Video S1201 S0303 S1301 S1701
Random $0.445$ $0.385$ $0.397$ $0.407$
Ours 0.867 0.735 0.760 0.713
@@ E.2 Comparison Experiments
We provide a comprehensive analysis of the comparative experimental results.
@@
Specifically, we provide detailed evaluations for within-scene (same scenes, same weather), cross-scene and cross-domain testing. Table 12 reports the $CR$ metric of three models under cross-scene and cross-domain conditions, while Table 13 presents the $TSR$ metric.
@@
As shown in Table 12 and Table 13, the proposed GC-VAT significantly outperforms SOTA methods. Due to the reward design based on physical distance, both the AOT e2e and D-VAT dvat fail to accurately reflect the agent’s tracking performance from a top-down perspective (see Appendix C.1 for theoretical proof), leading to misleading training signals for the tracker. Consequently, neither AOT nor D-VAT can effectively learn meaningful features, resulting in irregular performance distributions. In contrast, the proposed GC-VAT achieves superior convergence across all scenes. Specifically, in cross-scene experiments, the testing performance of the agent on the downtown map is relatively low, indicating that dense buildings and complex road elements pose significant challenges to the agent. Conversely, the testing performance on the village map is comparatively high, suggesting that the uniform color and simpler road conditions in the village map present fewer challenges.
@@
For cross-domain testing experiments, the agent performs well under night and foggy conditions but struggles under snow conditions. This indicates that the proposed GC-VAT exhibits strong robustness to changes in lighting and visibility but is less adaptive to variations in scene tone.